OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

2x-y=6 =>(•(-2)): -4x+2y=-12 4x-2y=12 4x-2y=12 -------------------- + 0=0 That it's true so the system of eq. has infinite solution; that means that the two lines are tangent (the one to the other).